When luke, the new kid, asked how old Annabelle was, she replied, "in two years, i'll be twice as old as i wa five years ago." How old is Annabelle?
1
Expert's answer
2011-09-06T11:10:02-0400
Let x be the age of Annabelle.
Then
x+2 = 2(x-5).
Hence
x+2 = 2x-10,
x=12